OpenStudy (anonymous):

how do you find the inverse of f(x)=x^5 and graph it?

OpenStudy (anonymous):

So basically set y = f(x), so you get: y = x^5. Then switch x and y, to get: x = y^5. Now solve it for y again, and that's the inverse function. You can then graph it like any other function.

OpenStudy (anonymous):

That's what i tried to do but my textbook says: f(x)=x^5 x^5=y x=sqrt 5 (y) y=sqrt 5 (x) f^-1(x)= sqrt 5(x) f^-1(x)= x^(1/5) How did the square root get there on the third equation?

OpenStudy (anonymous):

Basically it's just the fifth root. So you know how you do this: \[\sqrt[]{x^2} = x \rightarrow \sqrt[5]{x^5}=x\]

OpenStudy (anonymous):

Great Thank you~